OpenStudy (anonymous):

the polynomial 2x^3 + px^2 +qx +r is divisible by x^2-1, but when divided by x-3 it has remainder 8. what is (p-r).q

OpenStudy (anonymous):

a)10 b)-10 c)20 d)-20

OpenStudy (anonymous):

the first statement tells you two things if you replace \(x\) by 1 you get 0 and if you replace \(x\) by -1 you also get 0 last statement says if you replace \(x\) by 3 you get 8

OpenStudy (anonymous):

yes, i know.

OpenStudy (anonymous):

so unless there is some snappier way to do this, you have to solve the system of equations \[2+p+q+r=0\] \[-2+p-q+r=0\] \[54p+9p+3q+r=8\]

OpenStudy (anonymous):

check my arithmetic, first one i replaced x by 1 second one i replaces x by -1 and third one i replaced x by 3

OpenStudy (anonymous):

okay, thank you, i didnt know that i had to use systems of equations! i'll do that then! :)

OpenStudy (anonymous):

we can also write \[p+q+r=-2\] \[p-q+r=2\] \[9p+3q+r=-46\]

OpenStudy (anonymous):

i had a typo on the last one i put a p where there should not have been one

OpenStudy (anonymous):

its okay :) i'll try work it out later